Thinking about considerations of acid/base chemistry and solubility constants. Calcium carbonate (CaCO3) is added to initially pure water that is exposed to the atmosphere with 400 ppm(v) of carbon dioxide. If solid CaCO3 remains after a long time – showing dissolution and precipitation are at equilibrium – what is the aqueous concentration of calcium ions? Use the acid constants, Henry’s Law constants and solubility products.
Answer #1

Equilibrium of dissolution of CaCO3 is given by the equation :

CaCO3 (s) ⇌ Ca2+ (aq)+ CO32−(aq)

Ksp = 6×10−9

There are following equilibrium from CO2 in air to water for this:

1. CO2 (aq) + H2O    \rightleftharpoons H2CO3 (aq)

2. H2CO3 (aq) \rightleftharpoons H+ (aq) + HCO3- (aq) Ka1 =4.5*10-7

3. HCO3- (aq)  \rightleftharpoons H+ (aq) + CO32- (aq) Ka2 =4.7*10-11

Amount of CO2 Dissolved in be calculated from Henry 's law ;  

[CO2 (aq) ] = KCO2  * PCO2 (KCO2  = Henry constant for CO2 = 3.4 x 10-2 mol /(L-atm) )

Here we are using 'Henry solubility defined via concentration'.

At 400 ppm , PCO2 =4*10-4 atm

[CO2 (aq) ] =  3.4 x 10-2 mol /(L-atm) *4*10-4 atm = 1.36*10-5 mol/ L

[CO2 (aq) ] =  [ H2CO3 ]

so, from Ka, ; [HCO3-  ] = 2.5*10-6 mol/ L

and [CO32-  ]=1.1*10-8 mol/ L

we have,   [Ca2+  ]= [CO32-  ]= Ksp = 6×10−9

we have; [Ca2+  ]=s

and [CO32-  ]= (s + 1.1*10-8)

so, s* (s + 1.1*10-8) = 6×10−9

[Ca2+  ] = s = 7.8×10−5 M